# A Simple but Strong Baseline for Sounding Video Generation: Effective Adaptation of Audio and Video

### v1

@@ A.2 Implementation details
We used AnimateDiff (Guo et al., 2023) and AudioLDM (Liu et al., 2023a) as base models for video and audio, respectively. In both models, we insert the additional module just after each of the last two down-sampling blocks and each of the first two up-sampling blocks. Consequently, our model contains four additional modules at each modality. Following CoDi (Tang et al., 2023), each module is implemented by a single Transformer decoder block, while a cross-attention layer is replaced with a self-attention one for CMC-PE. We also followed the architecture of the environment encoder in CoDi for our connector. The total number of parameters in the newly added modules is about 468M. This is nearly four times smaller than that of the U-Nets in AnimateDiff and AudioLDM, which exceeds 1.7B.
@@
As mentioned previously, we basically optimize the newly added modules as well as the connectors while fixing the pre-trained parameters during training. However, we made one exception, namely, the motion layers in AnimateDiff, which are also fine-tuned during training. This is because these layers are dedicated to a certain frame rate and duration of videos (specifically, eight frames per second and two seconds), which do not necessarily match those of the training data (as shown in Section 4).
A.3 Captioning for GreatestHits dataset
We created video captions of the GreatestHits dataset utilizing LLaVA-NeXT (specifically LLaVA-NeXT-Video-7B-DPO released by Zhang et al. (2024b)). The text prompt is set as shown below:
What is being hit with a drumstick in the video? Please answer concisely with a brief description of its background by beginning with ”A person is hitting”.
A.4 Computation of AV-Align scores
A.4.1 Definition
AV-Align score Yariv et al. (2023) is defined as Intersection-over-Union (IoU) between onsets detected from the audio and peaks obtained from the optical flow. Specifically, it is computed as
[…]
where $\mathcal{A}$ represents a set of the onsets detected from the audio signal, and $\mathcal{V}$ represents a set of the peaks in the optical flow. A peak is considered to be valid in the other modality, if any corresponding peak exists within a window of three frames.
A.4.2 Official implementation
One issue in the computation of the AV-Align score is in evaluating $|\mathcal{A}\cup\mathcal{V}|$. As the temporal resolution of the audio is much higher than that of the video, a single peak in the video may have multiple corresponding peaks in the audio. In this case, there is no trivial way to count the number of elements in $\mathcal{A}\cup\mathcal{V}$ due to this one-to-many matching property. To avoid this problem, the official implementation adopts the following equation to compute the AV-Align score:
[…]
However, when $|\mathcal{A}|>|\mathcal{V}|$, the computed score can exceed one, which is unreasonable considering the original definition of the AV-Align score.
A.4.3 Modification
We modified the score computation so that it follows the original definition of the score. First, we rewrite IoU using precision and recall as
$\mathrm{IoU}=\frac{\mathrm{Precision}\cdot\mathrm{Recall}}{\mathrm{Precision}+\mathrm{Recall}-\mathrm{Precision}\cdot\mathrm{Recall}}.$ (13)
Utilizing this rewritten equation, we can compute the AV-Align score as
[…]
$\mathrm{where}\ p=\frac{1}{|\mathcal{A}|}\sum_{a\in\mathcal{A}}1[a\in\mathcal{V}],\ r=\frac{1}{|\mathcal{V}|}\sum_{v\in\mathcal{V}}1[v\in\mathcal{A}].$ (15)
By computing the score in this way, we can obtain a normalized value that is reasonable as IoU while avoiding the previously mentioned issue.

### v3

@@ III-E Implementation details
We used AnimateDiff [10] and AudioLDM [16] as base models for video and audio, respectively. In both models, we insert the additional module just after each of the last two down-sampling blocks and each of the first two up-sampling blocks. Consequently, our model contains four additional modules at each modality. Following CoDi [5], each module is implemented by a single Transformer decoder block, while a cross-attention layer is replaced with a self-attention one for CMC-PE. We also followed the architecture of the environment encoder in CoDi for our connector. The total number of parameters in the newly added modules is about 468M. This is nearly four times smaller than that of the U-Nets in AnimateDiff and AudioLDM, which exceeds 1.7B.
@@
As mentioned previously, we basically optimize the newly added modules as well as the connectors while fixing the pre-trained parameters during training. However, we made one exception, namely, the motion layers in AnimateDiff, which are also fine-tuned during training. This is because these layers are dedicated to a certain frame rate and duration of videos (specifically, eight frames per second and two seconds), which do not necessarily match those of the training data (as shown in Section IV).
IV Experiments
IV-A Experiments with a dedicated dataset for evaluating temporal alignment
IV-A1 Dataset and evaluation metrics
We extended the GreatestHits dataset [18] for our experiments. It contains 977 videos of humans hitting various objects with a drumstick in the scene. As the hitting sound and motion are dominant in the video, this dataset is suitable for evaluating the temporal alignment between the generated video and audio. We created video captions using LLaVA-NeXT [42] and utilized them as text conditions in our method.
We evaluated the quality of the generated data from three perspectives: video quality, audio quality, and temporal alignment. For the former two, we used FVD [43] and FAD [44], both of which are widely used in the literature. To measure how much the generated video and audio align with each other in terms of temporal dynamics, we used the AV-Align score proposed by [35]. This score is defined as Intersection-over-Union between onsets detected from the audio and peaks obtained from the optical flow. It is especially useful to measure the temporal alignment in the GreatestHits dataset, as hitting sounds make clear onsets, and hitting motions have correlating and distinct peaks in the optical flow.
We slightly modified how to compute AV-Align score from the official implementation. Specifically, we tuned hyper-parameters of the optical flow estimation and those of the onset detection to accurately estimate hitting timing using annotated timestamps in the Greatest Hits dataset. In addition, we computed the IoU after rewriting it with precision and recall to mitigate an issue caused by the difference in temporal resolution between video and audio. Details are provided in the Appendix VI-A.
IV-A2 Setup
  1. 1.
    One with the same setting as CoDi [5], in which cross-attention blocks are used for cross-modal conditioning.
  2. 2.
  3. 3.
For the training, we used the Adam optimizer [45] with a learning rate of 1e-5, and the batch size and the number of epochs were set to 16 and 1,000, respectively. For generation, we set the number of global timesteps $T$ to 25. We adopted classifier-free guidance [46] at each modality and set the strength of the guidance to 7.5 and 2.5 for video and audio, respectively, which are the standard settings in the original base models.
IV-A3 Results
Table I shows the evaluation results. Replacing cross-attention with CMC-PE improves the AV-Align score, which demonstrates that CMC-PE has a better inductive bias for temporal alignment than the cross-attention mechanism. The AV-Align score is further improved by conducting the timestep adjustment when generating data. This is achieved by making the generation process in both modalities mutually informative as discussed in Section III-B. Meanwhile, using too large a $\gamma$ leads to degradation of the performance due to the deviation from the original noise schedule. Overall, the proposed method performs substantially better in terms of the cross-modal alignment than our baseline following the design in CoDi [5].
On the other hand, while the alignment between audio and video has been improved by CMC-PE, it has come at the cost of a slight degradation of FVD. We conjecture that this is due to the inductive bias for higher temporal alignment induced by CMC-PE, which focuses more on boosting cross-modal alignment than bridging the domain gap between GreatestHits and the training data of AnimateDiff mentioned in Section III-E. A similar phenomenon has been also reported in the prior work [39].
TABLE I: Experimental results with the GreatestHits dataset.
Video quality Audio quality Alignment
Method FVD ($\downarrow$) FAD ($\downarrow$) AV-Align ($\uparrow$)
Cross-attention 379 2.35 0.250
CMC-PE 393 1.29 0.256
Ours ($\gamma=1.25$) 387 1.32 0.257
Ours ($\gamma=1.50$) 381 0.60 0.268
Ours ($\gamma=1.75$) 374 0.61 0.268
Ours ($\gamma=2.00$) 383 1.32 0.265
IV-B Experiments with benchmark datasets
IV-B1 Dataset and evaluation metrics
To compare the proposed method with existing methods, we conducted experiments with two popular benchmark datasets: Landscape [47] and VGGSound [48]. The Landscape dataset consists of 928 videos covering nine classes of natural scenes, while VGGSound is a substantially larger and more diverse dataset containing nearly 200K video clips covering about 300 sound classes. To enhance the data quality, we filtered 60K videos in which audio-video alignment is weak, as done in TempoToken [35]. In both datasets, we used the class names as the text conditions and trained our model to generate four-second audio-video pairs. The video comprises four frames per second, and the size of each frame is 256 $\times$ 256. The sampling rate of the audio is 16 kHz. Note that we changed fps from the previous experiments to follow the setting in the prior studies [23, 25, 35].
Differently from the previous experiments, we did not use AV-Align for the evaluation, as the videos in both Landscape and VGGSound often lack distinct motions highly correlating their audios. Instead, we used ImageBind score [49] between audio and video (IB-AV) to evaluate the cross-modal semantic alignment. Additionally, we computed the ImageBind score for text-audio and text-video pairs (IB-TA and IB-TV, respectively) to evaluate the audio and video quality in terms of fidelity to text condition.
IV-B2 Setup
  • T2A2V: We used TempoToken [35] to re-generate videos from the audios that are generated by our method.
  • T2V2A: We used SpecVQGAN [23] and DiffFoley [25] to re-generate audios from the videos that are generated by the proposed method.
  • Joint generation: We used MM-Diffusion [4]. For a fair comparison, the number of timesteps was set to be the same as that of the proposed method.
We chose these methods based on the availability of the official implementation provided by the authors. Note that the pretrained models of SpecVQGAN and DiffFoley were available for VGGSound, and that of MM-Diffusion was available for Landscape. When we cannot specify frame rate or resolution of the generated data, we resized the generated data to make it match our setting before the evaluation. For the training of our model, the batch size and the number of epochs were set to 16 and 100 for the Landscape dataset, and to 128 and 30 for VGGSound, respectively. The other settings are the same as those in the previous experiments.
IV-B3 Results
Tables II and III show the results with Landscape and VGGSound, respectively. Firstly, TempoToken failed to generate high-quality videos, which indicates that it is not robust against even small artifacts in the conditional audio caused by preceding text-to-audio generation. This is one of the most critical issues of sequential approaches, and SpecVQGAN and DiffFoley also suffered from it, resulting in relatively low audio quality. In contrast, the proposed method achieved the best quality in both video and audio except for IB-AV in VGGSound while attaining high cross-modal alignment. This indicates the importance of the training dedicated to joint generation and the effectiveness of our method.
Furthermore, the proposed model can be trained much more efficiently than the previous joint generation model. Our model only requires $\sim$2 hours with 8 A100 GPUs for training with Landscape dataset, whereas MM-Diffusion requires $\sim$60 hours with 32 A100 GPUs to train its base model, which generates videos with a resolution of 64x64. Such efficiency is achieved through the utilization of pre-trained diffusion models, making the advantage of our strategy clear.
TABLE II: Experimental results with Landscape dataset.
Video quality Audio quality Alignment
Method FVD $\downarrow$ IB-TV $\uparrow$ FAD $\downarrow$ IB-TA $\uparrow$ IB-AV $\uparrow$
TempoToken [35] $>$ 3000 0.220 0.146
MM-Diffusion [4] 1689 16.4 0.191
Proposed method 1122 0.238 6.63 0.146 0.192
TABLE III: Experimental results with VGGSound dataset. ($\dagger$ It uses a larger dataset for learning cross-modal alignment)
Video quality Audio quality Alignment
Method FVD $\downarrow$ IB-TV $\uparrow$ FAD $\downarrow$ IB-TA $\uparrow$ IB-AV $\uparrow$
TempoToken [35] 2473 0.155 0.168
SpecVQGAN [23] 5.08 0.059 0.100
DiffFoley [25] 5.72 0.074 0.159
Proposed method 333 0.277 1.46 0.129 0.155
V Conclusion